10. (Sections 11.1 - 11.3, 7.5, 9.3) (a) State the Implicit Function Theorem for an equation in the three variables x, y and z. (2) (b) Consider the equation xyz = 4x 2 + y 2 − z 2 . Use the Implicit Function Theorem to show that the given equation has a smooth unique local solution of the form z = g(x, y) about the point (2, 0, 4). Then find the local linearization of g about the point (2, 0). Hints: • Use the method of Example 11.2.6, but take into account that you are dealing with an equation in three variables here and that g in this case is a function of two variables. • Before you apply the Implicit Function Theorem, you should show that all the necessary conditions are satisfied. • Study Remark 11.3.3(1) and Remark 9.3.6(2).


1
Expert's answer
2021-04-28T10:42:55-0400

1.(a) The implicit function theorem addresses a question with two versions;

  • The analytic version implying a question about finding a solution of a system of non linear equations
  • The geometric version about the geometric structure of certain sets.

Assuming "s" is an open subset of "R^{n+k}" and that "F:S\to R^k" is a function of class "C^1" . Assume also that "(a,b)" is a point in that

"F(a,b)=0" and det "D_yF(a,b)\not =0"

Consider a continuously differentiable function "F(x,y,z)=c" such that

"dF\over dz" "(x,y,z)\not=0" then "F" is "\alpha(x,y,z)" such that "(x,y)" is sufficiently close to "(x,y)" then "F(x,y,z)=C"

10.(b) Given , "xyz=4x^2+y^2-z^2....(i)"

Let "F=4x^2+y^2-z^2-xyz"

"F=4x^2+y^2-z^2-xyz"

"\nabla F=[8x-yz,2y-xz,-2z-xy]"

"\nabla F(2,0,4)=[8\times 2-0\times 4,2\times 0-2\times 4, -2\times 4-2\times 0]"

"=[16, -8,-8]"

Equation of linearization

"z=z_o+F_x(x-xy)+F_y(y-y_1)....(ii)"

"z_o(x=2,y=0)" then "z=z_o"

From equation "(ii)"

"2(0).z_o=4\times (2)^2+0^2-z_o^2"

"2_o^2=16^2"

"z_o=4"

From equation"(ii)"

"z=4+16(x-2)-8(y-0)"

"=4+16x-32-8y,"

"16x-8y-z-28=0"